### v1

# View-Invariant Pixelwise Anomaly Detection in Multi-object Scenes with Adaptive View Synthesis

Abstract
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. Supervised segmentation methods can be applied to identify known problems, but unsupervised anomaly detection approaches are required when unknown anomalies occur. Current unsupervised pixel-level anomaly detection methods have mainly been developed for industrial settings where the camera position is known and constant. However, we find that these methods fail to generalize to the case when images are not perfectly aligned. We term the problem of unsupervised anomaly detection between two such imperfectly aligned sets of images as Scene Anomaly Detection (Scene AD). We present a novel network termed OmniAD to address the Scene AD problem posed. Specifically, we refine the anomaly detection method reverse distillation to achieve a 40% increase in pixel-level anomaly detection performance. The network’s performance is further demonstrated to improve with two new data augmentation strategies proposed that leverage novel view synthesis and camera localization to improve generalization. We validate our approach with qualitative and quantitative results on a new dataset, ToyCity, the first Scene AD dataset with multiple objects, as well as on the established single object-centric dataset, MAD. https://drags99.github.io/OmniAD/
Index Terms:
Anomaly Detection, Unsupervised Learning, Camera Localization
I Introduction
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (represented in green), and query views which may have anomalies in them (represented in red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous views and anomalous views.
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time [1, 2, 3, 4, 5]. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. For instance, the United States Navy places significant emphasis on the monitoring and detection of anomalies during inspections of their ships, which is critical for maintaining security and operational readiness and safety [6, 7]. Similarly, bridge inspectors can leverage this technology to identify parts of the bridge that are changing or behaving abnormally. This information is vital for predicting future performance and for maintenance decision making. In each of these examples, new image sets are acquired at every inspection but typically unaligned with previous inspections.
We define the challenge of detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views as Scene Anomaly Detection (Scene AD). The recaptured scene will be referred to as the query scene and the corresponding set of images as the query set. Importantly, the query scene may or may not contain any anomalies. Additionally, what may constitute as an anomalous region is not available in most scenarios, thus providing supervision for supervised methods is usually unfeasible. Therefore, Scene AD requires methods that do not rely on any supervision for what may constitute as an anomalous region.
Prior unsupervised anomaly detection methods have not been designed to handle Scene AD and are typically developed for controlled industrial settings with constant camera positions. As a result, we find that state-of-the art anomaly detection methods perform poorly on the Scene AD task. Apart from the architectures themselves, we find that there are two main challenges that need to be overcome. First, there is no benchmark multi-object dataset available for Scene AD. As described earlier, existing AD benchmarks typically focus on a single view. Zhou et al. proposed a multi-pose AD [8] dataset, however this problem setting focuses on multiple poses of a single object and not that of an entire scene. Second, any solution to Scene AD must thus incorporate a mechanism for view-invariance.
Due to the limited generalizability of prior unsupervised anomaly detection (AD) methods in Scene AD, we propose OmniAD, a refinement of Reverse Distillation (RD) [9]. To detect anomalous region locations using RD, deviations in a student decoder must occur with respect to a teacher encoder. We postulate that expanding the effective receptive field (ERF) [10] of the student decoder will increase the deviations produced by the student decoder, which in turn will increase the sensitivity to anomalous regions. To expand the ERF, we incorporate student attention modules into the student decoder. We show quantitatively that the expansion of the ERF will, in turn, increase the performance of RD.
However, to achieve improved view-invariance, we introduce two new data augmentation methods. Building on the intuition that it is much easier to compare perfectly aligned pictures, rather than comparing misaligned ones, our data augmentation methods utilize novel view synthesis (NVS) techniques. Specifically, we employ Neural Radiance Fields (NeRF) [11], combined with a camera localization strategy, hierarchical localization (hloc) [12, 13], to generate novel views of the query scene that closely align with the non-anomalous views. We hypothesize that providing these generated aligned novel views will allow AD methods to improve generalizability. Although supervised methods for anomaly detection cannot leverage this data augmentation due to the lack of ground truth for anomalies, unsupervised methods can effectively use this approach for improved performance.
To properly evaluate and benchmark the effectiveness of these methods, a suitable dataset is required. Recognizing the lack of a comprehensive benchmark for Scene AD in real scenes with multiple objects, we introduce the ToyCity dataset. ToyCity is tailored to provide a robust real benchmark featuring multiple camera viewpoints and a range of anomalies across different setups. ToyCity’s unique structure presents a new and significant challenge to state-of-the-art AD models. We evaluate several AD methods, including OmniAD, using ToyCity as well as an object focused dataset, the Multi-pose Anomaly Detection dataset (MAD) [8].
To summarize our main contributions are:
  • Introduction of the novel computer vision task of Scene AD involving detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views. Results are presented to demonstrate the inability of existing anomaly detection methods to generalize effectively to the Scene AD task.
  • The ToyCity dataset specifically designed for Scene AD, addressing the need for suitable benchmarks for multi-object and multi-view problems in this domain.
  • Development of a new anomaly detection architecture termed OmniAD that better generalizes to Scene AD. Specifically, we introduce a novel Student Attention Module to expand the effective receptive field (ERF) of the reverse distillation (RD) architecture and demonstrate state-of-the art performance on the ToyCity and MAD datasets [8].
  • Two novel data augmentation techniques leveraging Neural Radiance Fields and hierarchical localization, resulting in a 40% increase in performance for pixel-wise localization of anomalous regions in OmniAD.
This paper is organized as follows: II provides a review of the relevant literature in the area of anomaly detection, neural radiance fields, and hierarchical localization. III describes the proposed methodology for data augmentation as well as OmniAD. IV discusses the structure of the ToyCity dataset, developed to evaluate AD methods for scene AD. V describes the experimental setup, evaluation metrics, and both a quantitative and qualitative analysis of the experimental results. Finally, VI is a conclusion to the paper, summarizing the main findings, followed by references.
II Related Work
Anomaly Detection. AD algorithms typically harness a collection of non-anomalous images , to establish a normative distribution. AD is a critical function in industrial visual inspection for identifying defects [14] and is also employed in healthcare for the detection of irregularities in medical images [15]. Prominent AD methods include CFLOW-AD [16], FastFlow [17], CFA [18], DRÆM [19], and Reverse Distillation (RD) [9].
Both CFLOW-AD and FastFlow utilize a flow framework [20] that models the distribution of normal image patches to detect anomalous patterns. Typically, these methods begin by extracting features from a pretrained CNN encoder for an input image patch. The feature vectors are then mapped to a multivariate Gaussian distribution through the flow framework. Anomalous images can then be detected because they tend to fit poorly to this learned distribution, highlighting their abnormality.
The CFA method uses a combination of a frozen pretrained CNN, a patch descriptor network that learns on-the-fly, and a memory bank for storing compressed features. CFA fine-tunes the pretrained CNN features through metric learning, which is based on a coupled hypersphere approach, clustering normal features tightly within a hypersphere’s boundary.
DRÆM operates with two main networks: a reconstructive network and a discriminative network. DRÆM generates synthetic anomalies that the reconstructive network then transforms back to normal. The discriminative network uses both the reconstructed normal image and the synthetic anomalous image to predict segmentation for the regions where the synthetic anomaly was introduced.
RD adopts a knowledge distillation strategy [21, 22], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales. Unlike traditional approaches that feed the input image into both the student and the teacher networks, RD takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multi-scale feature map output by the teacher. This feature map is a compressed representation of the teacher’s features across multiple scales.
Predominant datasets for benchmarking these methods, like MVTecAD [14] and VisA [23], typically employ a fixed overhead camera perspective, resulting in a uniform view of anomalies. Our research demonstrates that significantly different views can lead to substantial degradation in the pixel-level localization of anomalies. To combat these limitations, our method leverages a novel view synthesis method and camera localization techniques to generate corresponding non-anomalous images, thus facilitating improved anomaly detection. While there are existing methods for 3D anomaly detection, they generally require 3D data that is aligned with a reference model.
Neural Radiance Field NeRFs, as introduced by Mildenhall et al. [11], represent a significant advancement in the field of view synthesis. These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene. Unlike conventional 3D Structure-from-Motion (SfM) methods that construct explicit geometric models, NeRFs model a scene’s volume through a neural network that predicts color and density for space points along camera rays. This technique enables the creation of photorealistic images from novel perspectives. Subsequent innovations like Nerfacto [24] and Instant-ngp [25] have further refined NeRFs, offering rapid training and real-time rendering capabilities. In the context of anomaly detection, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints, significantly improving the detection process.
Hierarchical Localization The process of hierarchical localization (hloc) [26, 13] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom (DoF) localization [26, 13, 27]. This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [12], Sift [28], or DISK [29]. A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [30]. The Perspective-n-Point algorithm, in conjunction with RANSAC, then utilizes these correspondences to estimate the camera’s pose. This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms. However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
III Method
An overview of our AD methodology for Scene AD can be seen in Figure2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
Figure 2: Overview of our methodology is depicted here. We utilize the non-anomalous views from a Scene AD to generate a non-anomalous SfM model. A novel view selection strategy utilizes anomaly views and the SfM model to generate novel views to augment the non-anomalous dataset. OmniAD is then trained on the augmented non-anomalous dataset for anomaly detection.
Figure 3: Overview of our novel view selection strategy. Utilizing the non-anomalous dataset, we generate a structure from motion (SfM) 3D model and accompanying camera pose for each image. Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation. Subsequently, we apply a NVS technique to generate images that present unique perspectives along the interpolated path. The anomaly dataset can also be used to determine more novel views.
III-A Novel View Selection Strategy
To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
III-A1 Interpolated Non-anomalous Views
Figure 4: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views. These interpolated views, not present in the original dataset, bridge the gap between the start and end poses, introducing new perspectives to the dataset.
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure4.
III-A2 Query-Aligned Non-anomalous Views
Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene. Our method employs hloc to accurately estimate the camera pose for a query anomaly image. This necessitates an accurate reconstruction of the baseline scene that is provided by the baseline SfM model. Hloc provides a variety of options for local feature detection, matching, and image similarity measurement. We opted for a combination of SuperPoint trained on Aachen and SuperGlue [12, 31], as it demonstrated effective performance on our datasets on average. Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.
III-B Novel View Synthesis
We selected Nerfacto [24] due to its balanced performance in training speed and quality compared to other methods, along with its user-friendly interface, which includes an interactive viewer. Although our method is agnostic to the specific NVS technique used, the aforementioned benefits made Nerfacto a compelling choice for this work. Continual advancements in this field, like the advent of 3D Gaussian Splattering[32], present opportunities for further enhancing rendering time and quality.
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure3. Example camera interpolated images can be seen in Figure3, and example images using QANV can be seen in Figure3
III-C OmniAD
We develop OmniAD to leverage our proposed augmentation techniques and surpass the performance of state-of-the-art AD methods when applied to Scene AD. OmniAD significantly enhances the RD framework through two significant updates, (i) integrating a ResNeXt backbone in the teacher encoders[33] (ii) and adding novel student attention modules, $A^{i}$. We found a combination of the ResNext with the student attention module to offer a performance increase of up to 40%, while the training and inference time cost for OmniAD is similar to RD.
Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder. We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
Student attention modules $A^{i}$ are introduced after each of the Student Decoder networks $D^{i}$ of the RD architecture. These modules are designed to bolster OmniAD’s sensitivity to anomalous samples. An illustrative overview of the OmniAD architecture is presented in Figure5. $A^{i}$ is based on the U-Net architecture and consists of alternating layers of 3x3 convolutions with self attention layers and a bottleneck layer consisting of several self attention layers as shown in Figure6. This architecture was designed to expand the Effective Receptive Field (ERF) of the Student Decoder which in turn would increase the variations in the output features in response to an anomaly. We validate this hypothesis and visualize the larger ERF in Figure7. This means there is a larger perturbation of features when an anomalous region is introduced into the model, making it more sensitive and responsive to such anomalies. This increased sensitivity is attributed to the improved performance of OmniAD. We conducted experiments to validate this hypothesis and demonstrate the efficacy of $A^{i}$.
Figure 6: Student attention module.
(a) with $A^{I}$
(b) w/o $A^{I}$
Figure 7: ERF of OmniAD with and without the student attention module. (a) with $A^{I}$. (b) w/o $A^{I}$.
IV ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure8. Each setup also contains different anomalies, as shown in Figure8. Additoinally, each anomaly image has at most one anomaly. The dataset includes the manually annotated ground truth for anomaly segmentation. Table I provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table II. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
Figure 8: ToyCity consists of four configurations as shown here. We present an example baseline image for each as well as an example of one of several anomalies used. A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
TABLE I: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
Name Non-Anomalous Anomalous QANV IBV
ToyCity 1 316 319 319 3781
ToyCity 2 341 308 308 3773
ToyCity 3 307 317 317 3673
ToyCity 4 314 314 314 3757
V Experiments
We evaluate several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
V-A Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$ (1)
The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous. This approach would yield numerous true negatives and few false positives. A similar effect can be observed with the AUROC, which may show a more optimistic estimate of model performance. However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives. Thus we provide both the $F_{1}$ score and AUROC.
V-B Implementation Details
To allow for fair comparison between all anomaly detection methods we use the anomalib [34] library implementation of each anomaly detection method. A fixed random seed of 42 was used for all methods. All methods were allowed to train for a maximum of 100 epochs and the max achieved $F_{1}$ score was selected as the best model. All methods converge well before 100 epochs were completed ensuring fair training. Additionally methods were trained with the same batch size. Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
V-C Ablation Studies
V-C1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables III and MAD Tables V we can see a consistent performance increase when $A^{i}$ is used.
V-C2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table II.
V-D Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables IV,VI. Additional experiments on different scenes is provided in the supplemental material.
The RD with ResNext model and the canonical RD methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model. The shift from v1 to v2, where anomaly aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies. Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ Scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations. This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and RD.
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table IV and MAD Table VI. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
TABLE II: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
Variation Non-Anomalous Anomalous QANV INV
v1 - -
v2 -
v3 -
v4
V-E Qualitative Evaluation
Figure9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
Figure 9: Segmentation performance comparison between OmniAD and RD with ResNeXt backbone across four ToyCity anomalies and 2 objects from MAD.We illustrating three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4m and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
Figure 10: Additional segmentations and anomaly scorces from OmniAD.
TABLE III: Comparison of Pixel $F_{1}$ scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
TABLE IV: Comparison of Pixel AUROC scores on ToyCity 1, 2, 3, and 4
Method ToyCity 1 ToyCity 2 ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
TABLE V: Comparison of Pixel $F_{1}$ scores on MAD-Real [8] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.672 0.714 0.680 0.635 0.334 0.395 0.471 0.524 0.341 0.368 0.374 0.328
OmniAD-R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
OmniAD-$A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
RD 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
CFLOW 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
FastFlow 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
CFA 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
TABLE VI: Comparison of Pixel AUROC scores on MAD-Real [8] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.999 0.999 0.999 0.999 0.988 0.993 0.999 0.999 0.941 0.968 0.973 0.971
OmniAD-R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
OmniAD-$A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
RD 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
CFLOW 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
FastFlow 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
CFA 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
VI Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in inspections of uncontrolled environments. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.