### v2

# OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis

Abstract
Visual anomaly detection in the built environment can be a valuable tool for applications such as infrastructure assessment, construction monitoring, security surveillance, and urban planning. Anomaly detection approaches are typically unsupervised and work by detecting deviations from an expected state where no assumptions are made exact type of deviation. Unsupervised pixel-level anomaly detection methods have been developed to successfully recognize and segment anomalies; however, existing techniques are designed for industrial settings with a fixed camera position. In the built environment, images may be periodically collected by a moving camera manually or by aerial or ground vehicles and utilized for decision making. The camera pose between successive collections may vary widely voiding a fundamental assumption in existing anomaly detection approaches. To address this gap, we introduce the problem of Scene Anomaly Detection (Scene AD), where the goal is to detect anomalies from two sets of images: one set without anomalies and one set that may or may not contain anomalies. No labeled semantic segmentation data are provided for training. We propose a novel network, OmniAD, to tackle Scene AD by refining the reverse distillation anomaly detection method, leading to a 40% improvement in pixel-level anomaly detection. Additionally, we introduce two new data augmentation strategies that leverage novel view synthesis and camera localization to enhance generalization. We evaluate our approach both qualitatively and quantitatively on a new dataset, ToyCity the first Scene AD dataset featuring multiple objects as well as on the established single object centric dataset, MAD. Our method demonstrates marked improvement over baseline approaches, paving the way for robust anomaly detection in scenes with real-world camera pose variations commonly observed in the built environment. Project Page: https://drags99.github.io/OmniAD/
keywords:
Anomaly Detection , Unsupervised Learning , Camera Localization , Neural Radiance Fields
[…]
[1] organization=Department of Electrical and Computer Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
[…]
[2] organization=Department of Civil and Environmental Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
1 Introduction
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (green), and query views which may contain anomalies (red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
For many applications in the built environment —ranging from infrastructure assessment and construction monitoring to security surveillance and urban planning—detecting visual anomalies in scenes captured over time is a critical task  [1, 2, 3, 4, 5]. These anomalies may arise from structural changes, newly introduced objects, unexpected modifications in appearance, or more generally, anything that deviates from an established non-anomalous reference. However, images acquired at different instances, whether manually or autonomously, typically suffer from misalignment due to variations in camera pose. For example, bridge inspectors may assess changes to a structure following a disaster, self-driving car systems might detect modifications to traffic signs, and the U.S. Military places strong emphasis on monitoring infrastructure for anomalies critical to operational readiness [6, 7, 8, 9, 10, 11, 12]. In all such cases, each new set of images is unaligned with previous captures, complicating anomaly identification.
We define the task of Scene Anomaly Detection (Scene AD) as detecting anomalous regions in a query set of images, potentially containing no anomalies at all, against a non-anomalous reference set. Labeled examples of what qualifies as an anomaly are unavailable, rendering supervised segmentation methods impractical. Additionally anomalies are often too infrequent or unpredictable to compile sufficient labeled training data for every potential anomaly type, rendering fully supervised models infeasible. As a result, unsupervised anomaly detection (AD) approaches are a feasible alternative. However, these AD methods were originally crafted for controlled industrial scenarios which commonly assume fixed camera views and struggle with the viewpoint variations of real-world Scene AD.
Two main challenges emerge. First, there is no suitable multi-object, multi-view benchmark for Scene AD: most existing AD datasets capture a single view or focus on a single object [13]. Second, any robust Scene AD solution must tolerate significant camera misalignment and still detect newly introduced or changed regions.
To address these gaps, we propose OmniAD, a refinement of the Reverse Distillation (RD) framework [14]. RD detects anomalies by comparing a student decoder’s reconstruction with a teacher encoder’s embedding. We hypothesize that broadening the student decoder’s effective receptive field (ERF) [15] will improve sensitivity to anomalous regions. Accordingly, we incorporate student attention modules to expand the ERF and enhance pixel-level anomaly detection.
Moreover, to manage viewpoint differences in real scenes, we introduce two data augmentation strategies leveraging Neural Radiance Fields (NeRF) [16] and hierarchical localization (hloc) [17, 18]. These techniques generate novel, more closely aligned views of the query scene, aiding unsupervised methods—such as OmniAD—in detecting anomalies without labeled examples.
Since no comprehensive benchmark exists for this multi-object, multi-view setting, we introduce the ToyCity dataset. ToyCity consists of scenes with multiple objects, varied camera viewpoints, and different anomalies. We evaluate OmniAD and several baseline methods on ToyCity as well as on the single-object-centric Multi-pose Anomaly Detection dataset (MAD) [13], demonstrating notable performance gains.
Our main contributions are:
  • 1.
    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery.
  • 2.
    ToyCity Dataset: A new multi-object, multi-view dataset designed for benchmarking Scene AD methods in realistic settings.
  • 3.
    OmniAD Architecture: A refined Reverse Distillation approach with a proposed Student Attention Module to expand the ERF, achieving state-of-the-art performance on ToyCity and MAD [13] and validated with ablation studies.
  • 4.
    NeRF-based Data Augmentation Methods: Two novel techniques using NeRF and hierarchical localization for improved view-invariance, yielding a 40% increase in anomaly localization accuracy.
The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization. Section 3 introduces the proposed OmniAD architecture and novel data augmentation strategies. Section 4 presents the ToyCity dataset. Section 5 details the experimental setup, metrics, and results. Section 6 discusses the limitations of our dataset and method. Finally, Section 7 concludes the paper.
2 Related Work
Anomaly Detection. AD algorithms typically harness a collection of non-anomalous images , to establish a normative distribution. AD is a critical function in industrial visual inspection for identifying defects [19] and is also employed in healthcare for the detection of irregularities in medical images [20]. Prominent AD methods include CFLOW-AD [21], FastFlow [22], CFA [23], DRÆM [24], and Reverse Distillation (RD) [14].
Both CFLOW-AD and FastFlow utilize a flow framework [25] that models the distribution of normal image patches to detect anomalous patterns. Typically, these methods begin by extracting features from a pretrained CNN encoder for an input image patch. The feature vectors are then mapped to a multivariate Gaussian distribution through the flow framework. Anomalous images can then be detected because they tend to fit poorly to this learned distribution, highlighting their abnormality.
The CFA method uses a combination of a frozen pretrained CNN, a patch descriptor network that learns on-the-fly, and a memory bank for storing compressed features. CFA fine-tunes the pretrained CNN features through metric learning, which is based on a coupled hypersphere approach, clustering normal features tightly within a hypersphere’s boundary.
DRÆM operates with two main networks: a reconstructive network and a discriminative network. DRÆM generates synthetic anomalies that the reconstructive network then transforms back to normal. The discriminative network uses both the reconstructed normal image and the synthetic anomalous image to predict segmentation for the regions where the synthetic anomaly was introduced.
RD adopts a knowledge distillation strategy [26, 27], where a student decoder network learns the feature space representation from a teacher encoder network at various feature map scales. Unlike traditional approaches that feed the input image into both the student and the teacher networks, RD takes an autoencoder-like approach where the teacher encodes the input image, and the student decodes a fused multi-scale feature map output by the teacher. This feature map is a compressed representation of the teacher’s features across multiple scales.
Predominant datasets for benchmarking these methods, like MVTecAD [19] and VisA [28], typically employ a fixed overhead camera perspective, resulting in a uniform view of anomalies. Our research demonstrates that significantly different views can lead to substantial degradation in the pixel-level localization of anomalies. To combat these limitations, our method leverages a novel view synthesis method and camera localization techniques to generate corresponding non-anomalous images, thus facilitating improved anomaly detection. While there are existing methods for 3D anomaly detection, they generally require 3D data that is aligned with a reference model.
Neural Radiance Field NeRFs, as introduced by Mildenhall et al. [16], represent a significant advancement in the field of view synthesis. These fields use a collection of images with known camera poses to synthesize new viewpoints of a scene. Unlike conventional 3D Structure-from-Motion (SfM) methods that construct explicit geometric models, NeRFs model a scene’s volume through a neural network that predicts color and density for space points along camera rays. This technique enables the creation of photorealistic images from novel perspectives. Subsequent innovations like Nerfacto [29] and Instant-ngp [30] have further refined NeRFs, offering rapid training and real-time rendering capabilities. In the context of anomaly detection, we leverage NeRFs to synthesize non-anomalous images from diverse viewpoints, significantly improving the detection process.
Hierarchical Localization The process of hierarchical localization (hloc) [31, 18] has been developed to accurately determine a camera’s position and orientation in space, a task known as 6-Degree-of-Freedom (DoF) localization [31, 18, 32]. This process begins with the extraction of local features from a query image, utilizing methods such as SuperPoint [17], Sift [33], or DISK [34]. A 3D SfM model serves as a reference for these local features, against which the query image is matched using global descriptors like VLAD [35]. The Perspective-n-Point algorithm, in conjunction with RANSAC, then utilizes these correspondences to estimate the camera’s pose. This pose is critical in enabling NeRFs to generate non-anomalous images that correspond to those containing anomalies, thus in theory providing a more robust dataset for anomaly detection algorithms. However as we shall show an improved AD method is needed to fully utilize these anomaly aligned images.
3 Method
An overview of our AD methodology for Scene AD can be seen in Figure  2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure  3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
Figure 2: Overview of our methodology is depicted here. We utilize the non-anomalous views from a Scene AD to generate a non-anomalous SfM model. A novel view selection strategy utilizes anomaly views and the SfM model to generate novel views to augment the non-anomalous dataset. OmniAD is then trained on the augmented non-anomalous dataset for anomaly detection.
Figure 3: Overview of our novel view selection strategy. Utilizing the non-anomalous dataset, we generate a structure from motion (SfM) 3D model and accompanying camera pose for each image. Leveraging these non-anomalous camera poses, we design a trajectory that connects them, allowing for interpolation. Subsequently, we apply a NVS technique to generate images that present unique perspectives along the interpolated path. The anomaly dataset can also be used to determine more novel views.
3.1 Novel View Selection Strategy
To generate new non-anomalous camera poses for NVS, we propose and evaluate two different approaches, namely Interpolated Non-anomalous Views (INV) and Query-aligned non-anomalous views (QANV).
3.1.1 Interpolated Non-anomalous Views
Figure 4: The provided sequence illustrates camera interpolation, where the pose transitions from the starting image to the ending image, are used to synthesize intermediate views. These interpolated views, not present in the original dataset, bridge the gap between the start and end poses, introducing new perspectives to the dataset.
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure  4.
3.1.2 Query-Aligned Non-anomalous Views
Here we propose a method to generate images of the baseline scene, but using poses from the anomaly scene. Our method employs hloc to accurately estimate the camera pose for a query anomaly image. This necessitates an accurate reconstruction of the baseline scene that is provided by the baseline SfM model. Hloc provides a variety of options for local feature detection, matching, and image similarity measurement. We opted for a combination of SuperPoint trained on Aachen and SuperGlue [17, 36], as it demonstrated effective performance on our datasets on average. Nonetheless, the flexibility to switch to other methods is inherent in our approach, enabling adaptation to specific dataset characteristics.
3.2 Novel View Synthesis
We selected Nerfacto [29] due to its balanced performance in training speed and quality compared to other methods, along with its user-friendly interface, which includes an interactive viewer. Although our method is agnostic to the specific NVS technique used, the aforementioned benefits made Nerfacto a compelling choice for this work. Continual advancements in this field, like the advent of 3D Gaussian Splattering[37], present opportunities for further enhancing rendering time and quality.
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure  3. Example camera interpolated images can be seen in Figure  3, and example images using QANV can be seen in Figure  3
3.3 OmniAD
Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level. Although Reverse Distillation (RD) [14] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
Backbone Exploration. We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints. Many alternatives struggled to converge or offered minimal gains, but ResNeXt [38] consistently yielded better multi-scale representations and improved performance compared to simpler encoders. In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder. However, given the ever-growing variety of such networks, we restricted our exploration to widely adopted backbones and found that ResNeXt provided the most consistent gains for our multi-view setting.
Receptive Field Constraints. Further analysis indicated that RD’s student decoder had a relatively small effective receptive field (ERF) [15], causing it to overlook anomalies. This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
Rationale of ERF Expansion for Anomaly Sensitivity. Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image. Suppose a teacher network $T$ extracts multi-scale features
$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$
where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map. A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$. In Reverse Distillation [14], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous (normal) data alone.
Effective Receptive Field (ERF). For each output location $(i,j)$ in $\hat{x}$, define its effective receptive field as
$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$
In standard RD, convolutions and limited skip connections yield a relatively localized ERF. If a true anomaly lies outside $\mathrm{ERF}_{ij}$ for many $(i,j)$, it may not sufficiently alter $\hat{x}$, causing the anomaly to go undetected.
Impact of ERF Expansion. Let $A\subset\mathcal{I}$ denote the anomaly region, i.e., pixels in $x$ that deviate from normal. By adding student attention modules (Fig. 6), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$:
$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$
Consequently, for any anomalous pixel $(u,v)\in A$, there are more $(i,j)$ in the reconstructed output $\hat{x}$ for which $(u,v)\in\mathrm{ERF}_{ij}(S^{\prime}\circ T)$. In other words, the anomaly influences a broader swath of the reconstruction, resulting in larger or more easily detectable errors.
Reverse Distillation Anomaly Score. As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring discrepancy between $\hat{x}_{ij}$ and $x_{ij}$. When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity. With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
Attention Modules for the Student Decoder. To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block. These modules expand the ERF by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies. Figure 6 shows the basic layout of each attention module, incorporating interleaved 3x3 convolutions and self-attention blocks for both local and global feature integration.
Figure 5: OmniAD’s architecture mirrors RD’s, featuring a frozen teacher encoder that channels features through a bottleneck module into a student decoder. We enhance this setup by integrating student attention modules $A^{1}$, $A^{2}$, and $A^{3}$, bolstering the architecture’s proficiency in utilizing Query-aligned non-anomalous views and Interpolated Baseline Views for improved effectiveness.
Figure 6: Student attention module.
(a) with $A^{I}$
(b) w/o $A^{I}$
Figure 7: ERF of OmniAD with and without the student attention module. (a) with $A^{I}$. (b) w/o $A^{I}$.
4 ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure  8. Each setup also contains different anomalies, as shown in Figure  8. Additionally, each anomaly image has at most one anomaly but each setup has varying anomalies that were added and removed. The dataset includes the manually annotated ground truth for anomaly segmentation. Table 1 provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table 2. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
Figure 8: ToyCity consists of four configurations as shown here. We present an example baseline image for each as well as an example of one of several anomalies used. A zoomed comparison is also presented that shows the anomaly highlighted with a red outline.
Table 1: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
Name Non-Anomalous Anomalous QANV IBV
ToyCity 1 316 319 319 3781
ToyCity 2 341 308 308 3773
ToyCity 3 307 317 317 3673
ToyCity 4 314 314 314 3757
5 Experiments
We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
5.1 Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.
$F_{1}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}$ (1)
The reason accuracy is deceptive is that a high accuracy metric can be achieved with a simplistic approach that labels all pixels as non-anomalous. This approach would yield numerous true negatives and few false positives. A similar effect can be observed with the AUROC, which may show a more optimistic estimate of model performance. However, it does not accurately reflect the model’s ability to precisely localize anomalies, primarily due to the abundance of true negatives. Thus we provide both the $F_{1}$ score and AUROC.
5.2 Implementation Details
To allow for fair comparison between all anomaly detection methods we use the anomalib [39] library implementation of each anomaly detection method. A fixed random seed of 42 was used for all methods. All methods were allowed to train for a maximum of 100 epochs and the max achieved $F_{1}$ score was selected as the best model. All methods converge well before 100 epochs were completed ensuring fair training. Additionally methods were trained with the same batch size. Other hyper-parameters were left as default values provided by anomalib for each anomaly detection method.
5.3 Ablation Studies
We conducted ablation studies to validate both the OmniAD architecture and the proposed NVS augmentation strategy.
5.3.1 OmniAD architecture
OmniAD has two major changes from RD. First we change the backbone to ResNeXt and second we introduced branching student attention modules $A^{i}$. To test if our changes are significant we test OmniAD with and without ResNeXt (OmniAD-R) and without $A^{i}$ (OmniAD-$A^{i}$) against RD. As shown in the $F_{1}$ scores on ToyCity Tables 3 and 4 and MAD Tables 7 we can see a consistent performance increase when $A^{i}$ is used.
5.3.2 NVS Augmentation
We test the efficacy of our NVS augmentation strategies by evaluating our metrics with and without each of the augmentations. Our hypothesis that generating novel views of the our non-anomalous scene similar to the view point of the query scene will make it easier for OmniAD to identify anomalies is verified through four different variations listed in Table 2.
5.4 Quantitative Evaluation
In our assessment of diverse AD methods on the ToyCity datasets, we observe a consistent pattern of performance increase OmniAD, OmniAD without ResNext (OmniAD w/o R), and RD when a novel view selection strategy is employed for $F_{1}$ scores on ToyCity and MAD. A similar trend can also be seen for AUROC on both ToyCity and MAD Tables 5,6,8. Additional experiments on different scenes is provided in the supplemental material.
The RD with ResNext model and the canonical RD methodology exhibit similar enhancement trends, yet they exhibit smaller gains in comparison to our model. The shift from v1 to v2, where anomaly aligned images are introduced, is particularly telling of the significant enhancements in the models’ ability to localize anomalies. Other AD methods such as CFLOW, FastFlow, CFA, and DRÆM show a range of responses to the augmentations, with $F_{1}$ Scores experiencing minimal improvements, remaining unchanged, or occasionally declining with the introduction of more complex augmentations. This suggests a potential lack of effective utilization of augmented data by these methods in contrast to OmniAD and RD.
As mentioned metrics such as pixel-level AUROC can often show very high numbers for datasets such as ToyCity and MAD due to the low number of pixels that may be an anomaly. This can be seen for both ToyCity Table 5,6 and MAD Table 8. Most AD methods will achieve an AUROC very close to 1 while still offering low $F_{1}$ scores.
Table 2: Four variations are tested for each of the scenes present in ToyCity. Variation v1 is un-augmented, v2 is with the Query-Aligned Views, v3 is with the interpolated Non-Anomalous views, and v4 is the combination of Query-Aligned Baseline views and Interpolated Non-Anomalous views.
Variation Non-Anomalous Anomalous QANV INV
v1 - -
v2 -
v3 -
v4
5.5 Qualitative Evaluation
Figure  9 illustrates the comparative segmentation capabilities of OmniAD and the second-best performing method, RD, which utilizes a ResNeXt backbone. A threshold was selected for each method that maximizes their pixel-level $F_{1}$ score. Three primary failure scenarios are evident. First, as depicted in the ToyCity 1 and 4 anomaly instances, RD may erroneously identify arbitrary areas as anomalous. Secondly, RD has a propensity to overlook entire anomalous regions, a shortcoming exemplified by the ToyCity 2 anomaly. Lastly, RD often overestimates the extent of an anomaly, erroneously marking larger areas than necessary. Similar defects in RD can also be seen from the MAD dataset in which it misses an anomaly or overestimates an anomalous region. Additional samples from OmniAD for each dataset is presented in Figure 10, showing a heat map of the anomaly score as well as the thresholded segmentation.
Figure 9: Segmentation performance comparison between OmniAD and RD with ResNeXt backbone across four ToyCity anomalies and 2 objects from MAD.We illustrating three primary failure patterns: misidentifying of non-anomalous areas (ToyCity 1,4m and MAD Pig), overlooking anomalies (ToyCity 2 and Mad Puppy), and overestimation of anomaly areas (ToyCity 3 and MAD Pig).
Figure 10: Additional segmentations and anomaly scorces from OmniAD.
Table 3: Comparison of Pixel $F_{1}$ scores on ToyCity 1 and 2
Method ToyCity 1 ToyCity 2
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.432 0.450 0.497 0.496 0.348 0.599 0.491 0.577
OmniAD-R 0.422 0.448 0.431 0.445 0.363 0.531 0.383 0.503
OmniAD-$A^{i}$ 0.391 0.411 0.409 0.388 0.256 0.546 0.476 0.357
RD 0.307 0.364 0.365 0.327 0.317 0.350 0.388 0.390
CFLOW 0.190 0.218 0.088 0.052 0.147 0.183 0.105 0.082
FastFlow 0.124 0.135 0.080 0.079 0.173 0.159 0.094 0.128
CFA 0.109 0.083 0.072 0.078 0.092 0.090 0.056 0.081
DRÆM 0.031 0.018 0.043 0.047 0.034 0.092 0.041 0.039
Table 4: Comparison of Pixel $F_{1}$ scores on ToyCity 3 and 4
Method ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.281 0.384 0.319 0.366 0.379 0.541 0.390 0.523
OmniAD-R 0.281 0.338 0.294 0.369 0.337 0.466 0.376 0.502
OmniAD-$A^{i}$ 0.196 0.276 0.243 0.290 0.279 0.434 0.284 0.403
RD 0.112 0.167 0.171 0.220 0.270 0.380 0.257 0.352
CFLOW 0.066 0.052 0.056 0.064 0.093 0.083 0.132 0.128
FastFlow 0.214 0.210 0.112 0.111 0.189 0.188 0.123 0.112
CFA 0.107 0.081 0.067 0.131 0.094 0.120 0.054 0.062
DRÆM 0.029 0.032 0.044 0.018 0.047 0.087 0.064 0.091
Table 5: Comparison of Pixel AUROC scores on ToyCity 1 and 2
Method ToyCity 1 ToyCity 2
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.990 0.996 0.999 0.999 0.991 0.998 0.999 0.999
OmniAD-R 0.988 0.993 0.999 0.999 0.991 0.995 0.997 0.999
OmniAD-$A^{i}$ 0.988 0.996 0.999 0.998 0.988 0.998 0.999 0.997
RD 0.987 0.992 0.998 0.998 0.986 0.990 0.998 0.998
CFLOW 0.920 0.914 0.864 0.921 0.928 0.934 0.935 0.868
FastFlow 0.912 0.939 0.940 0.956 0.948 0.952 0.957 0.952
CFA 0.943 0.946 0.976 0.976 0.953 0.962 0.977 0.984
DRÆM 0.578 0.449 0.585 0.597 0.621 0.600 0.700 0.725
Table 6: Comparison of Pixel AUROC scores on ToyCity 3 and 4
Method ToyCity 3 ToyCity 4
No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.979 0.993 0.997 0.999 0.982 0.993 0.995 0.998
OmniAD-R 0.980 0.985 0.998 0.998 0.980 0.989 0.996 0.998
OmniAD-$A^{i}$ 0.967 0.990 0.997 0.997 0.975 0.990 0.995 0.997
RD 0.961 0.987 0.995 0.996 0.977 0.990 0.995 0.997
CFLOW 0.914 0.844 0.906 0.914 0.932 0.930 0.899 0.916
FastFlow 0.946 0.932 0.947 0.936 0.949 0.947 0.971 0.972
CFA 0.928 0.953 0.975 0.979 0.928 0.953 0.975 0.979
DRÆM 0.616 0.628 0.612 0.626 0.627 0.570 0.594 0.626
Table 7: Comparison of Pixel $F_{1}$ scores on MAD-Real [13] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.672 0.714 0.680 0.635 0.334 0.395 0.471 0.524 0.341 0.368 0.374 0.328
OmniAD-R 0.613 0.677 0.621 0.623 0.302 0.346 0.454 0.441 0.312 0.204 0.297 0.284
OmniAD-$A^{i}$ 0.634 0.646 0.377 0.363 0.306 0.286 0.343 0.254 0.275 0.298 0.248 0.173
RD 0.616 0.628 0.520 0.199 0.219 0.203 0.302 0.260 0.262 0.236 0.170 0.192
CFLOW 0.496 0.466 0.484 0.579 0.162 0.142 0.150 0.147 0.309 0.163 0.214 0.209
FastFlow 0.646 0.651 0.575 0.598 0.066 0.087 0.043 0.058 0.297 0.285 0.282 0.297
CFA 0.568 0.576 0.594 0.067 0.294 0.232 0.048 0.029 0.189 0.151 0.117 0.139
DRÆM 0.042 0.016 0.022 0.018 0.040 0.041 0.072 0.110 0.020 0.041 0.037 0.028
Table 8: Comparison of Pixel AUROC scores on MAD-Real [13] of Puppy, Pig, and Elephant
Method Puppy Pig Elephant
No Aug QANV INV Both No Aug QANV INV Both No Aug QANV INV Both
OmniAD 0.999 0.999 0.999 0.999 0.988 0.993 0.999 0.999 0.941 0.968 0.973 0.971
OmniAD-R 0.999 0.993 0.999 0.999 0.988 0.983 0.999 0.999 0.924 0.892 0.933 0.928
OmniAD-$A^{i}$ 0.999 0.999 0.998 0.999 0.983 0.988 0.997 0.997 0.948 0.946 0.966 0.963
RD 0.999 0.999 0.999 0.999 0.981 0.982 0.998 0.998 0.936 0.929 0.951 0.956
CFLOW 0.996 0.996 0.999 0.999 0.962 0.967 0.993 0.990 0.945 0.906 0.947 0.953
FastFlow 0.999 0.998 0.999 0.999 0.868 0.896 0.965 0.966 0.862 0.914 0.934 0.932
CFA 0.998 0.998 0.999 0.994 0.977 0.976 0.986 0.979 0.916 0.927 0.934 0.942
DRÆM 0.819 0.359 0.514 0.524 0.720 0.571 0.684 0.796 0.601 0.496 0.559 0.519
6 Limitations and Future Work
OmniAD and the proposed ToyCity dataset address unsupervised Scene Anomaly Detection in multi-view scenarios, yet several limitations remain. First, each ToyCity image contains at most one anomaly at a time, even though multiple anomalies were added and removed from each scene. In contrast, real-world settings such as disaster assessments or industrial inspections may involve multiple anomalies within a single capture. Note that ToyCity does include partial occlusions, and OmniAD itself does not strictly limit detection to a single anomaly region. Second, although ToyCity was captured with real cameras at varying angles and lighting conditions, its tabletop-scale scenes do not fully encompass the complexities of large scale infrastructure.
Furthermore, our data augmentation assumes sufficient coverage for camera localization. In highly variable environments or when dealing with reflective surfaces, generating high-quality synthetic views can be challenging without accurate camera localization. Nonetheless, as shown in our experiments, OmniAD’s performance does not hinge solely on these augmented views.
In future work, we plan to:
  • 1.
    Extend ToyCity to include multiple anomalies in a single image and introduce dynamic elements such as moving vehicles or pedestrians.
  • 2.
    Validate OmniAD on larger real-world datasets (e.g., complete bridge or ship inspections) to further assess its robustness.
  • 3.
    Investigate faster 3D reconstruction, localization, and noval view synthesis methods.
  • 4.
    Identifying the underlying failure modes in GAN/autoencoder based AD methods under significant viewpoint variation remains an open research question, which we leave for future work.
7 Conclusion
Our study introduced and addressed a specific anomaly detection (AD) problem we term Scene AD. Scene AD deals with pixel-level AD between a non-anomalous scene and an anomalous scene, where a misalignment between non-anomalous and anomalous views is expected, as is often the case in applications in the built environment. To challenge and measure the effectiveness of state-of-the-art anomaly detection for Scene AD, we introduced the ToyCity dataset. This new benchmark focuses on viewpoint variance in scenes of multiple objects and serves as a testbed for our new method OmniAD. We demonstrated through extensive comparisons that OmniAD outperforms other AD methods when applied to the ToyCity dataset as well as the object centric established MAD-Real, establishing a new state-of-the-art for this task. Particularly, we found that the incorporation of a student attention module increases the efficacy of reverse distillation for pixel-level AD by expanding the effective receptive filed (ERF). Our experiments also demonstrated the benefits of novel view synthesis and camera localization for dataset augmentation in enhancing the robustness and accuracy of OmniAD allowing for an increase in pixel level $F_{1}$ of up to 40% in many scenes. We hope that our work will catalyze further advancements, spurring the creation of advanced anomaly detection techniques capable of managing the complexities associated with varying viewpoints in 3D scenes.
8 Acknowledgments
The authors acknowledge partial financial support from the Texas Department of Transportation (Grant No. 0-7181), and Department of Defense (Project No. G0511607). The contents of this paper reflect the views of the authors, who are responsible for the facts and the accuracy of the data presented herein. The contents do not necessarily reflect the official views or policies of the sponsors. The authors acknowledge the use of the Carya Cluster and the advanced support from the Research Computing Data Core at the University of Houston to carry out the research presented here.